Jason takes off from rest across level water on his jet-powered skis. The combined mass of Jason and his skis is 75 kg(the mass of the fuel is negligible). The skis have a thrust of 200N and a coefficient of kinetic friction on water of 0.10. Unfortunately, the skis run out of fuel after only 67s. How far has Jason traveled when he finally coasts to a stop?

A) 7600 m

B) 3800 m

C) 10,000 m

D) 5400 m
Answer #1

net acceleration of jason during thrust

a1 = F/m - u g

a1 = 200/75 - 0.1* 9.8

a1 = 1.687 m/s^2

velocity after 67 s

v = a1 * t = 1.687* 67 = 113 m/s

distance traveled during 67 s

d1 = 0.5 a1* t^2 = 0.5* 1.687* 67^2

d1 = 3786.47 m

now, acceleration after thrust stopped

a2 = 0.1* 9.8 = 0.98 m/s^2

distance traveled during this time

d2 = v^2/ ( 2 a2)

d2 = 113^2 / ( 2* 0.98)

d2 = 6514.8 m

total distance traveled

d = d1 + d2 = 10000 m

====

Comment before rate in case any doubt, will reply for sure.. goodluck